OpenStudy (anonymous):

G is the incenter of Δ ABC. m∠A = 55 m∠B = 70 m∠C = 55 What is the m∠ABF?

OpenStudy (anonymous):

help me on that

OpenStudy (anonymous):

umm okay

OpenStudy (anonymous):

wait on secc

OpenStudy (anonymous):

lmao wow

OpenStudy (anonymous):

A number is divided by 6. Then, 3 is added to the quotient, and the result is 5. What is the number? A.(4)/(3) B. 12 C. 18 D. 48

OpenStudy (anonymous):

The term "A number" refers to any number, obviously, so usually we use the letter "X" So, it's asking what is x / 6 + 3 = 5 First, lets get rid of the +3 in the equation. x / 6 + 3 = 5 -3 -3 Now, the problem is x / 6 is positive 2 So, multiply 6 to to both sides to remove the fraction x / 6 = 2 *6 *6 x = 12

OpenStudy (anonymous):

Sooo your answer is B